# Beyond The Rainbow: High Performance Deep Reinforcement Learning On A Desktop PC

### v1

@@ Appendix G BTR with and without Epsilon Greedy
One of the first observations we made early in the testing process was that the inclusion of using $\epsilon$-greedy in addition to NoisyNetworks benefited some environments but not others. Specifically, performance was reduced on BattleZone and Phoenix, both games where the agent reached very high levels of performance with extremely precise control. However, DoubleDunk performed significantly worse, only reaching a score of 0, rather than the score of 23 the final BTR algorithm achieved. Similar findings were also found in the full version of Rainbow DQN which used only NoisyNetworks, which achieved a best score of -0.3 (Dopamine’s “compact” Rainbow DQN however which did not use NoisyNetworks achieved 22). From this, we conclude that NoisyNetworks alone failed to sufficiently explore the environment, whereas $\epsilon$-greedy did not. From these results, we eventually decided to use both methods, but disable $\epsilon$-greedy halfway through training to reap the best of both techniques.
Appendix H Experiment Compute Resources
H.1 Our Compute Resources
@@
When testing ideas originally (those mentioned in Appendix I), we only tested them using a single run of the games BattleZone, NameThisGame and Phoenix unless otherwise stated. Whilst this method of evaluation is not statistically significant, for preliminary purposes with computational restrictions, we deemed this the best option.
H.2 BTR with Different Hardware
In this work, we look to make high-performance RL more accessible to those with less compute resources, especially those only with access to desktop computers. Most of our experiments were performed with an RTX4090, we also provide some walltimes for 200M Atari frames for lower-end machines, and provide a brief comparison of desktop PCs against internal clusters:
@@ H.2 BTR with Different Hardware
Appendix I Other Things We Tried
@@
Recently in RL, it has been shown that Neural Networks have a severe problem with under-parameterisation (Kumar et al., 2020) and neurons becoming dormant (Sokar et al., 2023), preventing larger models from seeing the success evident in other areas of Deep Learning. One method used to remedy this is weight decay through the use of the AdamW optimizer (Loshchilov & Hutter, 2017). As this is far simpler than many other the other techniques to prevent under-parameterisation we decided to test this approach, however no significant differences in performance were observed. We tested this approach using the decay parameter $41e-$ 4, however potentially using a higher value may results in significant changes in performance.
@@
Shortly after the submission of this work, we tested BTR with addition of Layer Normalization, and found positive results. Layer Normalization can improve the robustness to a variety of pathologies that cause loss of plasticity (Lyle et al., 2024), and helps to improve the conditioning of the network’s gradients in RL (Ball et al., 2023). Below in Figures I7, I8 and Table I7, we show the results of this addition into BTR.
Figure I7: Graph shows individual game performance of BTR with and without Layer Normalization. Layer Normalization makes a notable improvement on multiple games, including NameThisGame and BattleZone.
Figure I8: Graph shows IQM human-normalized performance of BTR with and without Layer Normalization on the Atari-5 Benchmark.
Table I7: Maximum scores obtained during training (averaged over 100 episodes and all performed random seeds) after 200M Frames on the Atari-5 Environment, compared to BTR with Layer Normalization.
@@
NameThisGame 2292 8049 32834 28710
Phoenix 761 7243 498264 367284
QBert 164 13455 48485 45034
IQM 0.000 1.000 8.369 7.627
Median 0.000 1.000 5.801 4.589
Mean 0.000 1.000 21.099 16.561
Appendix J Altered Atari Environment Settings
In order to investigate the impact of the environmental sticky actions parameter and to compare against other works, we include results for it on the Atari-5 benchmark in Figure J9.
Figure J9: Graph shows individual game performance of BTR with and without Sticky Actions. NameThisGame and Phoenix see small improvements. When using sticky actions, the impact of disabling $\epsilon$-actions is far more noticeable ($\epsilon$-actions are disabled after 125M frames). This indicates using sticky actions produces a policy more robust to noise, as would be expected.
Some prior works choose to pass life information to the agent (Schmidt & Schmied, 2021). To clarify, this is different to terminal on loss of life. Life information does not reset the episode upon losing a life, but does pass a terminal to the buffer, allowing the agent to experience further into episodes while also giving the agent a negative signal for losing a life. This setting is not recommended in Machado et al. (2018), and works which use it are not comparable to those which don’t. To emphasize this point, we take the three games from the Atari-5 benchmark which use lives (as not all Atari games do), and perform a comparison.
Figure J10: BTR with and without life information.
Appendix K BTR for Wii Games
BTR interfaces with different Wii Games via the Dolphin Emulator. Specifically, we use a forked repository of to allow Python scripts to interact with the emulator. This includes loading savestates (used to reset episodes), grabbing the screen as a PIL (Clark, 2015) image at the Wii’s internal resolution of 480p (downsampled to 140x114 and grey-scaled, used for all observations), reading the Wii’s RAM (used for reward functions and termination conditions) and allows programmatic input into the emulator (used for actions). Using Dolphin’s portable setting, we are able to run multiple Dolphin Emulators simultaneously on the same machine. Each instance runs as a unique process, and communicates with the agent via Python’s multiprocessing library. Similarly to the Atari benchmark, for all games we used a frameskip of 4.
K.1 Super Mario Galaxy
@@
K.2 Mario Kart Wii
The Mario Kart Wii environment had the agent play against the game’s internal opponents (on hard mode, with 12 racers including the agent), on the course Rainbow Road (with items on the 150cc speed setting). The agent had to complete 3 laps of the course to finish the race. The agent had just 4 action, including accelerate, drifting left or right, and using its item. While this limited the agent’s potential actions substantially, we found using fewer actions to dramatically accelerate training.

### v2

@@ Appendix F BTR with and without Epsilon Greedy
One of the first observations we made early in the testing process was that the inclusion of using $\epsilon$-greedy in addition to NoisyNetworks benefited some environments but not others. Specifically, performance was reduced on BattleZone and Phoenix, both games where the agent reached very high levels of performance with extremely precise control. However, DoubleDunk performed significantly worse, only reaching a score of 0, rather than the score of 23 the final BTR algorithm achieved. Similar findings were also found in the full version of Rainbow DQN, which used only NoisyNetworks, which achieved a best score of -0.3 (Dopamine’s “compact” Rainbow DQN, however, which did not use NoisyNetworks achieved 22). From this, we conclude that NoisyNetworks alone failed to sufficiently explore the environment, whereas $\epsilon$-greedy did not. From these results, we eventually decided to use both methods, but disable $\epsilon$-greedy halfway through training to reap the best of both techniques.
Appendix G Experiment Compute Resources
G.1 Our Compute Resources
@@
When testing ideas originally (those mentioned in Appendix H), we only tested them using a single run of the games BattleZone, NameThisGame and Phoenix unless otherwise stated. Whilst this method of evaluation is not statistically significant, for preliminary purposes with computational restrictions, we deemed this the best option.
G.2 BTR with Different Hardware
In this work, we look to make high-performance RL more accessible to those with fewer computing resources, especially those only with access to desktop computers. Most of our experiments were performed with an RTX4090, we also provide some walltimes for 200M Atari frames for lower-end machines and provide a brief comparison of desktop PCs against internal clusters:
@@ G.2 BTR with Different Hardware
Appendix H Other Things We Tried
@@
  • Using Exponential Moving Average networks rather than using fixed target networks (this was both computationally slower and performed worse).
  • Varying the frequency of updating the target network (we tested 250, 500 and 1000, finding 500 to perform best).
  • Changing the size of maxpool layer following the convolutional layers (we tested 4 and 8, however 6 performed significantly better).
@@
[…]
  • Using the AdamW optimizer(Loshchilov & Hutter, 2019) which uses weight decay with the decay parameter $41e-$ 4, however found this made no significant difference.
  • Using the GeLu activation instead of ReLu, which drastically reduced performance.
Only testing on a single environment (BattleZone), we also tried:
  • Annealing the discount rate from 0.97 to 0.997 throughout training, but found no significant difference.
  • Applying spectral normalization to the linear layers (dramatically worse performance).
  • Increasing the number of cos’ from IQN (no significant difference on performance).
  • Using Dopamine’s Prioritized Experience Replay buffer which doesn’t include a $\alpha$ value (moderately worse performance).
  • As discussed in F, we also tried not using $\epsilon$-greedy when using noisy nets.
@@
Shortly after the submission of this work, we tested BTR with addition of Layer Normalization, and found positive results. Layer Normalization can improve the robustness to a variety of pathologies that cause loss of plasticity (Lyle et al., 2024), and helps to improve the conditioning of the network’s gradients in RL (Ball et al., 2023). Below in Figure H9 we show the impact of including layer normalization into BTR.
Figure H9: Graph shows Atari-5 performance with and without layer normalization using Inter-quartile mean and Atari-60 predicted median from Aitchison et al. (2023). Layer normalization uses 3 seeds, with shaded areas showing 95% confidence intervals.
Table H9: Maximum scores obtained during training (averaged over 100 episodes and all performed random seeds) after 200M Frames on the Atari-5 Environment, compared to BTR with Layer Normalization.
@@
NameThisGame 2292 8049 33258 27917
Phoenix 761 7243 493762 427481
QBert 164 13455 47384 42927
IQM 0.000 1.000 8.191 7.739
Median 0.000 1.000 5.379 4.766
Mean 0.000 1.000 20.836 18.453
Appendix I Altered Atari Environment Settings
In order to investigate the impact of the environmental sticky actions parameter and to compare against other works, we include results for it on the Atari-5 benchmark in Figure I10.
Figure I10: Graph shows Atari-5 performance with and without sticky actions (sticky actions is the default) using Inter-quartile mean and Atari-60 predicted median from Aitchison et al. (2023). No Sticky Actions uses a single seed, so this result should be used with caution.
Some prior works choose to pass life information to the agent (Schmidt & Schmied, 2021). To clarify, this is different to terminal on loss of life. Life information does not reset the episode upon losing a life, but does pass a terminal to the buffer, allowing the agent to experience further into episodes while also giving the agent a negative signal for losing a life. This setting is not recommended in Machado et al. (2018), and works which use it are not comparable to those which don’t. To emphasize this point, we take the three games from the Atari-5 and perform a comparison.
Figure I11: Graph shows Atari-5 performance with and without life information using Inter-quartile mean and Atari-60 predicted median from Aitchison et al. (2023). Life Information uses 3 seeds, with shaded areas showing 95% confidence intervals. From this we conclude results using life information are invalid for comparison.
Appendix J BTR for Wii Games
BTR interfaces with different Wii Games via the Dolphin Emulator. Specifically, we use a forked repository of Dolphin Emulator to allow Python scripts to interact with the emulator. This includes loading savestates (used to reset episodes), grabbing the screen as a PIL image at the Wii’s internal resolution of 480p (downsampled to 140x114 and grey-scaled, used for all observations), reading the Wii’s RAM (used for reward functions and termination conditions) and allowed programmatic input into the emulator (for setting actions). Using Dolphin’s portable setting, we are able to run multiple Dolphin Emulators simultaneously on the same machine. Each instance runs as a unique process, and communicates with the agent via Python’s multiprocessing library. Similarly to the Atari benchmark, for all games we used a frameskip of 4.
J.1 Super Mario Galaxy
@@ Appendix I Altered Atari Environment Settings
J.2 Mario Kart Wii
The Mario Kart Wii environment had the agent play against the game’s internal opponents (on hard mode, with 12 racers including the agent), on the course Rainbow Road (with items on the 150cc speed setting). The agent had to complete 4 laps of the course to finish the race. The agent had just 4 action, including accelerate, drifting left or right, and using its item. While this limited the agent’s potential actions substantially, we found using fewer actions to dramatically accelerate training.
@@
Appendix K Atari-5 Regression Procedure
In our main paper ablation figure (Figure 5), we used the regression procedure recommended in Atari-5 (Aitchison et al., 2023). This procedure is typically used to predict the Median score across the entire 57 game Atari suite, while only needing to use 5 games. While we believe this procedure produces a valid and useful plot, we find that BTR did differ significantly from the predicted value. We opted to use both the predicted median and the IQM across the 5 games to give two easy to interpret averages. Figure K12 shows the 57 game suite’s true median, compared to the median predicted by Atari-5.
Figure K12: BTR’s 60 game median against that predicted by Atari-5 using the regression procedure from Aitchison et al. (2023). Shaded areas show 95% confidence intervals using 4 seeds.